OpenStudy (anonymous):

Solve the following by substitution: y = x 5x = 12y

OpenStudy (earthcitizen):

it's a false expression

OpenStudy (anonymous):

why?

OpenStudy (earthcitizen):

coz you get 5y=12y which is false or just zero

OpenStudy (anonymous):

I said it was 0,0 and I got it wrong

OpenStudy (earthcitizen):

hmm, try using Cramer's rule

OpenStudy (earthcitizen):

lets say x-y=0 5x-12y=0 then\[\left[\begin{matrix}1 & -1 \\ 5 & -12\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}0 \\ 0\end{matrix}\right)\]

OpenStudy (anonymous):

dam nikga nd i got it wrong lol.......

OpenStudy (earthcitizen):

Ax=b\[x=A ^{-1}*b\]

OpenStudy (anonymous):

my teachz dum

OpenStudy (earthcitizen):

lol

OpenStudy (earthcitizen):

\[\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}0 \\ 0\end{matrix}\right)\]